# Model-agnostic Origin Attribution of Generated Images with Few-shot Examples

### v1

@@ Appendix 0.L Model Attribution with Multiple Source Models
Since this involves multi-source origin attribution, the evaluation metric used in this section is accuracy instead of AUC. We evaluate the model under the following conditions: 2 classes (SD, VQ-D), 4 classes (SD, VQ-D, LDM, Glide), 6 classes (SD, VQ-D, LDM, Glide, GALIP, ProGAN), and 8 classes (SD, VQ-D, LDM, Glide, GALIP, ProGAN, StyleGAN2, GauGAN). Table 28 shows the accuracy of verifying multiple sources using the eight trained one-class classifiers. Table 29 shows the accuracy of directly training a multi-class classifier.
Appendix 0.M Comparision with Related Work
There are two other related works. However, they have different settings and use methods distinct from ours.
[53] proposed a method called RONAN (Reverse-engineering-based Origin Attribution) for origin attribution. The authors aimed to create an alteration-free method that does not impair generation quality and is suitable for pre-trained models. RONAN is based on a reverse-engineering algorithm applied to a source model, utilizing the reconstruction loss of reverse-engineering for inference. This process requires access to the source model and generates many additional images for auxiliary determination. For instance, to determine if an image is generated by a specific model, 100 images need to be generated for inference via Grubbs’ Hypothesis Testing [12].
[22] proposed a method called FLIPAD (Final-Layer Inversion Plus Anomaly Detection) to determine if a sample is generated by a specific model. The authors aimed to develop a method not limited to a closed-world setting and without the need for undesirable changes to the target model. FLIPAD is based on final-layer inversion for feature extraction and anomaly detection for inference. This process requires access to the source model and needs 10,000 samples during training.
In contrast, we aim to conduct origin attribution in an open world with only access to a few images generated by a source model. We formulate this problem as a few-shot one-class classification task. Our method, OCC-CLIP, is based on the pre-trained CLIP model and utilizes prompt engineering and adversarial data augmentation. It does not require access to any model and does not need to generate extra images for auxiliary determination. In the training phase, we only need 50, or even 10, samples from both non-target and target image datasets to perform effectively.
@@ A FOUNTAIN IN A ROUNDABOUT WITH PEOPLE PASSING BY
Table 10: Evaluation of OCC-CLIP, along with 14 basic methods. Training samples are sourced from SD (denoted as target) and COCO (denoted as non-target). The test data are assembled from SD as well as a different generative image dataset. The optimal outcomes for individual datasets are emphasized using bold formatting.
@@
Table 11: Evaluation of OCC-CLIP, along with 14 basic methods. Training samples are sourced from SD (denoted as target) and COCO (denoted as non-target). The test data are assembled from SD as well as a different generative image dataset. The optimal outcomes for individual datasets are emphasized using bold formatting.
@@
Table 28: Employing an ensemble of one-class classifiers for multi-class classification tasks. The following scenarios are considered: 2 classes, 4 classes, 6 classes, and 8 classes. The standard deviations of all the values are less than 0.04.
@@
8 classes CoOp $0.7334$
OCC-CLIP $\mathbf{0.7678}$
Table 29: Directly training a multi-class classifier based on the OCC-CLIP or CoOp. The following scenarios are considered: 2 classes, 4 classes, 6 classes, and 8 classes. The standard deviations of all the values are less than 0.04.

### v2

@@ Appendix 0.L Model Attribution with Multiple Source Models
Since this involves multi-source origin attribution, the evaluation metric used in this section is accuracy instead of AUC. We evaluate the model under the following conditions: 2 classes (SD, VQ-D), 4 classes (SD, VQ-D, LDM, Glide), and 6 classes (SD, VQ-D, LDM, Glide, GALIP, ProGAN). Table 31 shows the accuracy of verifying multiple sources using the eight trained one-class classifiers. Table 32 shows the accuracy of directly training a multi-class classifier.
Appendix 0.M Stronger Baselines and Harder Datasets
We also experimented with three representative methods of fine-tuning CLIP using LoRA: replacing the linear layer in the last (LoRAl), middle (LoRAm), or all (LoRAa) residual attention blocks. The other two baseline methods are Wang et al.’s  [64], which uses inverse engineering and has model access, and Girish et al.’s  [13], which uses a multistep pipeline. Table 29 shows the accuracy of these methods. Our approach outperforms these baselines in our setting.
@@ A FOUNTAIN IN A ROUNDABOUT WITH PEOPLE PASSING BY
Table 10: Evaluation of OCC-CLIP, along with 14 basic methods. Training samples are sourced from SD (denoted as target) and COCO (denoted as non-target). The test data are assembled from SD as well as a different generative image dataset. The optimal outcomes for individual datasets are emphasized using bold formatting. The metric shown here is AUC.
@@
Table 11: Evaluation of OCC-CLIP, along with 14 basic methods. Training samples are sourced from SD (denoted as target) and COCO (denoted as non-target). The test data are assembled from SD as well as a different generative image dataset. The optimal outcomes for individual datasets are emphasized using bold formatting. The metric shown here is AUC.
@@
Table 12: Evaluation of OCC-CLIP, along with 14 basic methods. Training samples are sourced from SD (denoted as target) and COCO (denoted as non-target). The test data are assembled from SD as well as a different generative image dataset. The optimal outcomes for individual datasets are emphasized using bold formatting. The metric shown here is Accuracy.
Methods VQ-D LDM Glide GALIP
VGG16 [55] $-20.6024\scriptscriptstyle\pm\scriptstyle 2.78e\text{-}$ 2 $-20.5260\scriptscriptstyle\pm\scriptstyle 1.53e\text{-}$ 2 $-20.5396\scriptscriptstyle\pm\scriptstyle 1.63e\text{-}$ 2 $-20.6501\scriptscriptstyle\pm\scriptstyle 3.44e\text{-}$ 2
ResNet50 [20] $-20.6235\scriptscriptstyle\pm\scriptstyle 4.71e\text{-}$ 2 $-20.5328\scriptscriptstyle\pm\scriptstyle 2.49e\text{-}$ 2 $-20.5764\scriptscriptstyle\pm\scriptstyle 3.03e\text{-}$ 2 $-20.6729\scriptscriptstyle\pm\scriptstyle 3.43e\text{-}$ 2
Inception-v3 [58] $-20.5999\scriptscriptstyle\pm\scriptstyle 3.32e\text{-}$ 2 $-20.5193\scriptscriptstyle\pm\scriptstyle 1.29e\text{-}$ 2 $-20.5219\scriptscriptstyle\pm\scriptstyle 1.31e\text{-}$ 2 $-20.6297\scriptscriptstyle\pm\scriptstyle 2.99e\text{-}$ 2
DenseNet-121 [22] $-20.6548\scriptscriptstyle\pm\scriptstyle 3.33e\text{-}$ 2 $-20.5198\scriptscriptstyle\pm\scriptstyle 2.07e\text{-}$ 2 $-20.5423\scriptscriptstyle\pm\scriptstyle 4.13e\text{-}$ 2 $-20.6731\scriptscriptstyle\pm\scriptstyle 4.53e\text{-}$ 2
ViT-B-16 [11] $-20.6849\scriptscriptstyle\pm\scriptstyle 2.16e\text{-}$ 2 $-20.5188\scriptscriptstyle\pm\scriptstyle 1.45e\text{-}$ 2 $-20.5633\scriptscriptstyle\pm\scriptstyle 2.55e\text{-}$ 2 $-20.7113\scriptscriptstyle\pm\scriptstyle 1.58e\text{-}$ 2
DeiT-B-16 [61] $-20.6285\scriptscriptstyle\pm\scriptstyle 3.15e\text{-}$ 2 $-20.5245\scriptscriptstyle\pm\scriptstyle 1.41e\text{-}$ 2 $-20.5401\scriptscriptstyle\pm\scriptstyle 2.25e\text{-}$ 2 $-20.6119\scriptscriptstyle\pm\scriptstyle 2.07e\text{-}$ 2
CaiT-S-24 [62] $-20.6070\scriptscriptstyle\pm\scriptstyle 3.82e\text{-}$ 2 $-20.5216\scriptscriptstyle\pm\scriptstyle 1.54e\text{-}$ 2 $-20.5394\scriptscriptstyle\pm\scriptstyle 3.41e\text{-}$ 2 $-20.6027\scriptscriptstyle\pm\scriptstyle 3.71e\text{-}$ 2
Swin-B-4 [36] $-20.7796\scriptscriptstyle\pm\scriptstyle 3.39e\text{-}$ 2 $-20.6498\scriptscriptstyle\pm\scriptstyle 3.57e\text{-}$ 2 $-20.6652\scriptscriptstyle\pm\scriptstyle 2.93e\text{-}$ 2 $-20.7623\scriptscriptstyle\pm\scriptstyle 3.86e\text{-}$ 2
Image_Patch [39] $-20.6143\scriptscriptstyle\pm\scriptstyle 2.78e\text{-}$ 2 $-20.5136\scriptscriptstyle\pm\scriptstyle 1.13e\text{-}$ 2 $-20.5348\scriptscriptstyle\pm\scriptstyle 2.06e\text{-}$ 2 $-20.6341\scriptscriptstyle\pm\scriptstyle 4.01e\text{-}$ 2
Feature_Patch [39] $-20.6434\scriptscriptstyle\pm\scriptstyle 1.97e\text{-}$ 2 $-20.5767\scriptscriptstyle\pm\scriptstyle 2.81e\text{-}$ 2 $-20.6719\scriptscriptstyle\pm\scriptstyle 5.89e\text{-}$ 2 $-20.5908\scriptscriptstyle\pm\scriptstyle 3.95e\text{-}$ 2
CLIP [44] $0.5800\scriptscriptstyle\pm\scriptstyle 0.0$ $0.5695\scriptscriptstyle\pm\scriptstyle 0.0$ $0.5475\scriptscriptstyle\pm\scriptstyle 0.0$ $0.5755\scriptscriptstyle\pm\scriptstyle 0.0$
CoOp [71] $-20.8457\scriptscriptstyle\pm\scriptstyle 5.91e\text{-}$ 2 $-20.6745\scriptscriptstyle\pm\scriptstyle 7.54e\text{-}$ 2 $-20.8041\scriptscriptstyle\pm\scriptstyle 7.07e\text{-}$ 2 $-20.8722\scriptscriptstyle\pm\scriptstyle 5.88e\text{-}$ 2
OCC-CLIP $\mathbf{0.9118\scriptscriptstyle\pm\scriptstyle 1.73e\text{-}2}$ $\mathbf{0.7654\scriptscriptstyle\pm\scriptstyle 3.28e\text{-}2}$ $\mathbf{0.8783\scriptscriptstyle\pm\scriptstyle 2.20e\text{-}2}$ $\mathbf{0.9227\scriptscriptstyle\pm\scriptstyle 1.91e\text{-}2}$
Table 13: Evaluation of OCC-CLIP, along with 14 basic methods. Training samples are sourced from SD (denoted as target) and COCO (denoted as non-target). The test data are assembled from SD as well as a different generative image dataset. The optimal outcomes for individual datasets are emphasized using bold formatting. The metric shown here is Accuracy.
Methods ProGan StyleGan2 GauGan Overall
VGG16 [55] $-20.6218\scriptscriptstyle\pm\scriptstyle 4.14e\text{-}$ 2 $-20.5939\scriptscriptstyle\pm\scriptstyle 5.57e\text{-}$ 2 $-20.6597\scriptscriptstyle\pm\scriptstyle 3.82e\text{-}$ 2 $-20.5991\scriptscriptstyle\pm\scriptstyle 3.53e\text{-}$ 2
ResNet50 [20] $-20.6157\scriptscriptstyle\pm\scriptstyle 3.92e\text{-}$ 2 $-20.5910\scriptscriptstyle\pm\scriptstyle 6.70e\text{-}$ 2 $-20.6858\scriptscriptstyle\pm\scriptstyle 2.48e\text{-}$ 2 $-20.6140\scriptscriptstyle\pm\scriptstyle 4.07e\text{-}$ 2
Inception-v3 [58] $-20.6045\scriptscriptstyle\pm\scriptstyle 3.65e\text{-}$ 2 $-20.5711\scriptscriptstyle\pm\scriptstyle 3.59e\text{-}$ 2 $-20.6455\scriptscriptstyle\pm\scriptstyle 2.96e\text{-}$ 2 $-20.5846\scriptscriptstyle\pm\scriptstyle 2.88e\text{-}$ 2
DenseNet-121 [22] $-20.6556\scriptscriptstyle\pm\scriptstyle 3.34e\text{-}$ 2 $-20.5967\scriptscriptstyle\pm\scriptstyle 5.24e\text{-}$ 2 $-20.7276\scriptscriptstyle\pm\scriptstyle 2.66e\text{-}$ 2 $-20.6243\scriptscriptstyle\pm\scriptstyle 3.76e\text{-}$ 2
ViT-B-16 [11] $-20.5649\scriptscriptstyle\pm\scriptstyle 2.79e\text{-}$ 2 $-20.6148\scriptscriptstyle\pm\scriptstyle 4.20e\text{-}$ 2 $-20.6606\scriptscriptstyle\pm\scriptstyle 2.62e\text{-}$ 2 $-20.6169\scriptscriptstyle\pm\scriptstyle 2.62e\text{-}$ 2
DeiT-B-16 [61] $-20.5451\scriptscriptstyle\pm\scriptstyle 2.04e\text{-}$ 2 $-20.5601\scriptscriptstyle\pm\scriptstyle 3.47e\text{-}$ 2 $-20.6051\scriptscriptstyle\pm\scriptstyle 2.08e\text{-}$ 2 $-20.5736\scriptscriptstyle\pm\scriptstyle 2.44e\text{-}$ 2
CaiT-S-24 [62] $-20.5437\scriptscriptstyle\pm\scriptstyle 2.33e\text{-}$ 2 $-20.5468\scriptscriptstyle\pm\scriptstyle 3.63e\text{-}$ 2 $-20.6166\scriptscriptstyle\pm\scriptstyle 2.68e\text{-}$ 2 $-20.5683\scriptscriptstyle\pm\scriptstyle 3.12e\text{-}$ 2
Swin-B-4 [36] $-20.5378\scriptscriptstyle\pm\scriptstyle 1.82e\text{-}$ 2 $-20.6057\scriptscriptstyle\pm\scriptstyle 3.40e\text{-}$ 2 $-20.6929\scriptscriptstyle\pm\scriptstyle 3.75e\text{-}$ 2 $-20.6705\scriptscriptstyle\pm\scriptstyle 3.31e\text{-}$ 2
Image_Patch [39] $-20.6128\scriptscriptstyle\pm\scriptstyle 4.33e\text{-}$ 2 $-20.6230\scriptscriptstyle\pm\scriptstyle 7.13e\text{-}$ 2 $-20.7010\scriptscriptstyle\pm\scriptstyle 3.19e\text{-}$ 2 $-20.6048\scriptscriptstyle\pm\scriptstyle 3.95e\text{-}$ 2
Feature_Patch [39] $-20.6582\scriptscriptstyle\pm\scriptstyle 2.29e\text{-}$ 2 $-20.7678\scriptscriptstyle\pm\scriptstyle 1.65e\text{-}$ 2 $-20.7333\scriptscriptstyle\pm\scriptstyle 2.30e\text{-}$ 2 $-20.6632\scriptscriptstyle\pm\scriptstyle 3.28e\text{-}$ 2
CLIP [44] $0.5415\scriptscriptstyle\pm\scriptstyle 0.0$ $0.5860\scriptscriptstyle\pm\scriptstyle 0.0$ $0.5170\scriptscriptstyle\pm\scriptstyle 0.0$ $0.5645\scriptscriptstyle\pm\scriptstyle 0.0$
CoOp [71] $-20.7053\scriptscriptstyle\pm\scriptstyle 8.57e\text{-}$ 2 $-20.8540\scriptscriptstyle\pm\scriptstyle 5.87e\text{-}$ 2 $-20.8854\scriptscriptstyle\pm\scriptstyle 3.84e\text{-}$ 2 $-20.8059\scriptscriptstyle\pm\scriptstyle 6.54e\text{-}$ 2
OCC-CLIP $\mathbf{0.8635\scriptscriptstyle\pm\scriptstyle 5.26e\text{-}2}$ $\mathbf{0.8939\scriptscriptstyle\pm\scriptstyle 2.69e\text{-}2}$ $\mathbf{0.9409\scriptscriptstyle\pm\scriptstyle 1.79e\text{-}2}$ $\mathbf{0.8824\scriptscriptstyle\pm\scriptstyle 2.94e\text{-}2}$
@@
Table 29: Additional Baselines (tested on VQ-D, LDM, Glide, GALIP, ProGAN, StyleGAN2, GauGAN). The table shows the overall average accuracy tested on 7 different testing tasks.
Methods ViT-L ViT-Res CLIP-LoRAm CLIP-LoRAl
Average ACC $0.5966$ $0.6269$ $0.5646$ $0.5507$
Methods CLIP-LoRAa Girish et al. Wang et al. OCC-CLIP
Average ACC $0.5951$ $0.6079$ $0.7600$ $\mathbf{0.8824}$
@@
Table 31: Employing an ensemble of one-class classifiers for multi-class classification tasks. The following scenarios are considered: 2 classes, 4 classes, and 6 classes. The standard deviations of all the values are less than 0.04.
@@
Table 32: Directly training a multi-class classifier based on the OCC-CLIP or CoOp. The following scenarios are considered: 2 classes, 4 classes, and 6 classes. The standard deviations of all the values are less than 0.04.